Right angle triangle. Hypotenuse is 7".
Aspect ratio is 16:9 (1.78:1)
a^2 * b^2 = c^2    
a=1.78x   b=x   c=7
(1.78x)^2 * x^2 = 7^2
(1.78x)^2 * x^2 = 49

Answer by checkley77(12844)

16:9 (1.78:1)
a^2 * b^2 = c^2
a=1.78x b=x c=7
[NOTE: YOU SHOULD BE ADDING, NOT MULTIPLYING, THE A & B TERMS]
(1.78x)^2 + x^2 = 7^2
(1.78x)^2 + x^2 = 49
3.1684x^2+x^2=49
4.1684x^2=49
x^2=49/4.1684
x^2=11.755
X=SQRT11.755
X=3.43
PROOF:
(1.78*3.43)^2+3.43^2=49
6.1^2+11.77=49
37.21+11.77=49
49~49